OpenStudy (anonymous):

f(x)=x+5 and g(x)=x^2. Find (g . f)(4)

OpenStudy (anonymous):

\(f\circ g\) or \(f\times g\) ?

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

start by getting rid of the circle notation and write \[f\circ g(x)=f(g(x))\] you should always start that way

OpenStudy (anonymous):

then replace the general \(g\) by the specific one you have, in this case it is \[g(x)=x^2\] so you get \(f(x^2)\)

OpenStudy (anonymous):

since \(f(x)=x+5\) you get \(f(x^2)=x^2+5\)

OpenStudy (anonymous):

but i just looked above and in your problem you only need \(f\circ g(4)\) so first find \(g(4)=4^2=16\)

OpenStudy (anonymous):

then take \(f(16)=16+5=21\)

OpenStudy (anonymous):

notice that you will get the same answer if you replace \(x\) by \(4\) in the function \(x^2+5\) either way

OpenStudy (anonymous):

yeah